OpenStudy (anonymous):

A business consultant earns a flat fee for his work as well as an hourly fee. He charges his clients at a rate of $75 per hour. In one week, he billed a client $1,800 for 20 hours worked. Which of the following equations describes the pay for the consultant in terms of hours worked?

OpenStudy (anonymous):

A. y = 20x – 300 B. y = 20x + 300 C. y = 75x – 300 D. y = 75x + 300

OpenStudy (debbieg):

Since x=hours worked, and he charges 75/hour for the hourly portion of the charge, you can easily eliminate 2 of the four choices, right? The hourly portion of the charge is 75(# of hours worked) and # of hours worked is represented by the variable x. Now he also charged a fixed portion, that does not depend on hours worked. So would that be an amount that is ADDED to the hourly portion, or would it be SUBTRACTED from the hourly portion? Get it? :)